# Scalable Valuation of Human Feedback through Provably Robust Model Alignment

### v1

@@ G.2 IPO do not satisfy the redescending property
~$\widetilde{\mathcal{L}}_{\mathrm{rDPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\coloneqq\frac{(1-c)\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]-c\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(-g_{\theta}(\tilde{s}))]}{1-2c},$ (32)
where $20\leq c<1/$ 2.
@@
$\mathrm{IF}_{\mathrm{rDPO}}(x,\theta,p_{\mathcal{D}})=\bigg{|}\bigg{(}\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{rDPO})}(s)\right]\bigg{)}^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{rDPO})}(s_{\textrm{flip}})]\bigg{|},$ (33)
@@
$\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{rDPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\beta\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg{[}\xi_{\theta^{*}(\epsilon)}(\tilde{s})\bigg{(}\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{win}}\mid\tilde{x})-\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{lose}}\mid\tilde{x})\bigg{)}\bigg{]},$
@@
$\xi_{\theta^{*}(\epsilon)}(\tilde{s})\coloneqq\frac{1-c}{1-2c}\sigma(-g_{\theta^{*}(\epsilon)}(\tilde{s}))+\frac{c}{1-2c}\sigma(g_{\theta^{*}(\epsilon)}(\tilde{s})).$
@@
00 0 $=\frac{\partial}{\partial\epsilon}\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{rDPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\bigg{|}_{\theta=\theta^{*}(\epsilon)}$
@@
$=-\beta\bigg{\{}\int\bigg{\{}\frac{\partial}{\partial\epsilon}p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\bigg{\}}F_{\theta^{*}(\epsilon)}^{(\textrm{rDPO})}(\tilde{s})\mathrm{d}\tilde{s}+\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg{[}\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}H_{\theta^{*}(\epsilon)}^{(\textrm{rDPO})}(\tilde{s})\bigg{]}\bigg{\}},$ (34)
@@
($\bigg{(}\int\bigg{\{}\frac{\partial}{\partial\epsilon}p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\bigg{\}}F_{\theta^{*}(\epsilon)}^{(\textrm{rDPO})}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}\bigg{|}_{\epsilon=0}=\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}(\epsilon)}^{(\textrm{rDPO})}(s_{\textrm{flip}})],$
@@
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Let $20\leq c<1/$ 2. Then, under Theorem 8, the IF of rDPO do not satisfy the robustness condition in Definition 2, i.e., $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{rDPO}}(x,\theta,p_{\mathcal{D}})\neq 0$.
@@
From the positive definite assumption in Theorem 8 and the $L$-Lipschitsness assumption of $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ in $\theta$, we can see that $F_{\theta^{*}}^{(\textrm{rDPO})}(s)$ is $(\xi_{\theta^{*}}(s)\cdot L)$-Lipschits continuous in $\theta$. We thus have $0<\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{rDPO})}(s)]\leq\xi_{\theta^{*}}(s)\cdot L<\infty$ under $20\leq c<1/$ 2, where we use Lemma 1.2.2 in Nesterov [75].
@@
Furthermore, from the assumption that $\|\nabla_{\theta}\log\pi_{\theta}(y\mid x)\|\leq C$ ($0<C<\infty$), we have $\|\nabla_{\theta}\log\pi_{\theta}(y_{\text{win}}\mid x)-\nabla_{\theta}\log\pi_{\theta}(y_{\text{lose}}\mid x)\|\leq 2C$ from the triangle inequality. Thus, the term $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{rDPO}}(x,\theta,p_{\mathcal{D}})$ can be upper bounded as $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{rDPO}}(x,\theta,p_{\mathcal{D}})\leq 2C(1-c)/L^{\prime}(1-2c)$, where $0<L^{\prime}\leq\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}(s)]$ and we use the fact that $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\xi_{\theta^{*}}(s_{\textrm{flip}})=\frac{1-c}{1-2c}$. The fact $0<2C(1-c)/L^{\prime}(1-2c)<\infty$ according to $0<C<\infty$, $20\leq c<1/$ 2, and $0<L^{\prime}<\infty$ completes the proof. ∎

### v2

@@ Appendix G IF Analysis for the DPO variants (summarized in Theorem 3)
$\widetilde{\mathcal{L}}_{\mathrm{rDPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\coloneqq\frac{(1-c)\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]-c\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(-g_{\theta}(\tilde{s}))]}{1-2c},$ (28)
where $0\leq c<1/2$.
@@
$\mathrm{IF}_{\mathrm{rDPO}}(x,\theta,p_{\mathcal{D}})=-\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{rDPO})}(s)\right]\bigg)^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{rDPO})}(s_{\textrm{flip}})],$ (29)
@@
$)],\displaystyle\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{rDPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\beta\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg[\xi_{\theta^{*}(\epsilon)}(\tilde{s})\bigg(\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{win}}\mid\tilde{x})-\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{lose}}\mid\tilde{x}$$\bigg)\bigg$],
@@
$).\displaystyle\xi_{\theta^{*}(\epsilon)}(\tilde{s})\coloneqq\frac{1-c}{1-2c}\sigma(-g_{\theta^{*}(\epsilon)}(\tilde{s}))+\frac{c}{1-2c}\sigma(g_{\theta^{*}(\epsilon)}(\tilde{s}$)).
@@
$0$ $=\frac{\partial}{\partial\epsilon}\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{rDPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\bigg|_{\theta=\theta^{*}(\epsilon)}$
@@
[…]
@@
$\bigg(\int\bigg\{\frac{\partial}{\partial\epsilon}p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\bigg\}F_{\theta^{*}(\epsilon)}^{(\textrm{rDPO})}(\tilde{s})\mathrm{d}\tilde{s}\bigg)\bigg|_{\epsilon=0}=\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}(\epsilon)}^{(\textrm{rDPO})}(s_{\textrm{flip}})],$
@@
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Let $0\leq c<1/2$. Then, under Theorem 6, the IF of rDPO do not satisfy the robustness condition in Definition 2, i.e., $\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{rDPO}}(x,\theta,p_{\mathcal{D}})\|\neq 0$.
@@
From Theorem 6, the IF for rDPO is given by
$\mathrm{IF}_{\mathrm{rDPO}}=-\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{rDPO})}(s)\right]\bigg)^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}(\epsilon)}^{(\textrm{rDPO})}(s_{\textrm{flip}})].$
From the positive definite assumption on the Hessian $H_{\theta^{*}}^{(\textrm{rDPO})}(s)$, it follows that its expectation $\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{rDPO})}(s)]$ is also a positive definite matrix. Let $L^{\prime}=\lambda_{\mathrm{min}}(\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{rDPO})}(s)])>0$ be its minimum eigenvalue. Then the norm of its inverse is bounded: $\|(\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{rDPO})}(s)])^{-1}\|\leq 1/L^{\prime}$.
Furthermore, from the assumption that $\|\nabla_{\theta}\log\pi_{\theta}(y\mid x)\|\leq C$, we have $\|\nabla_{\theta}\log\pi_{\theta}(y_{\text{win}}^{\textrm{flip}}\mid x)-\nabla_{\theta}\log\pi_{\theta}(y_{\text{lose}}^{\textrm{flip}}\mid x)\|\leq 2C$.
Taking the limit required by Definition 2, we have:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}$ $\|\mathrm{IF}_{\mathrm{rDPO}}\|$
$\leq\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{rDPO})}(s)\right]\bigg)^{-1}\bigg\|\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}(\epsilon)}^{(\textrm{rDPO})}(s_{\textrm{flip}})]\bigg\|$
$\leq(1/L^{\prime})\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}\bigg[\underbrace{\xi_{\theta^{*}}(s_{\textrm{flip}})}_{\text{IF Weight}}\cdot\underbrace{\|\nabla_{\theta}\log\pi_{\theta}(y_{\text{win}}^{\textrm{flip}}\mid x)-\nabla_{\theta}\log\pi_{\theta}(y_{\text{lose}}^{\textrm{flip}}\mid x)\|}_{\leq 2C}\bigg]$
$\leq(1/L^{\prime})\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}[\xi_{\theta^{*}}(s_{\textrm{flip}})]\cdot 2C,$
where
$\xi_{\theta^{*}}(s_{\textrm{flip}})=\frac{1-c}{1-2c}\sigma(-g_{\theta}(s_{\mathrm{flip}}))+\frac{c}{1-2c}\sigma(g_{\theta}(s_{\mathrm{flip}})).$
We now evaluate the limit of the IF weight $\xi_{\theta^{*}}(s_{\textrm{flip}})$:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\xi_{\theta^{*}}(s_{\textrm{flip}})=\left(\frac{1-c}{1-2c}\cdot 1\right)+\left(\frac{c}{1-2c}\cdot 0\right)=\frac{1-c}{1-2c}.$
By the bounded convergence theorem, the limit of the expectation is the expectation of the limit. Thus, the IF limit is upper bounded by:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{rDPO}}\|\leq(1/L^{\prime})\cdot\left(\frac{1-c}{1-2c}\right)\cdot 2C=\frac{2C(1-c)}{L^{\prime}(1-2c)}.$
The fact $0<\frac{2C(1-c)}{L^{\prime}(1-2c)}<\infty$ (according to $0<C<\infty$, $0\leq c<1/2$, and $0<L^{\prime}<\infty$) completes the proof. ∎
G.2 cDPO do not satisfy the redescending property
The objective of cDPO [76] is as follows:
$)].\displaystyle\widetilde{\mathcal{L}}_{\mathrm{cDPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\coloneqq(1-c)\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]-c\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(-g_{\theta}(\tilde{s}$))]. (31)
We first show the IF for the cDPO.
Theorem 7.
$\mathrm{IF}_{\mathrm{cDPO}}(x,\theta,p_{\mathcal{D}})=-\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{cDPO})}(s)\right]\bigg)^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{cDPO})}(s_{\textrm{flip}})],$ (32)
Proof.
The proof follows from the same argument as in Theorem 6, ignoring the (12c)(1-2c) term in the denominator. ∎
Corollary 4 (cDPO is not robust).
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Let $0\leq c<1$. Then, under Theorem 7, the IF of cDPO do not satisfy the robustness condition in Definition 2, i.e., $\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{cDPO}}(x,\theta,p_{\mathcal{D}})\|\neq 0$.
Proof.
@@
By following the proof in Corollary 3 and ignoring the (12c)(1-2c) term in the denominator, we have $0<\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{cDPO})}(s)]\leq\xi_{\theta^{*}}(s)\cdot L<\infty$ and thus $\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{cDPO}}(x,\theta,p_{\mathcal{D}})\|_{2}\leq 2C(1-c)/L^{\prime}$, where $0<L^{\prime}\leq\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}(s)]$ and we use the fact that $\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\xi_{\theta^{*}}(s_{\textrm{flip}})=1-c$. The fact $0<2C(1-c)/L^{\prime}<\infty$ according to $0<C<\infty$, $0\leq c<1$, and $0<L^{\prime}<\infty$ completes the proof. ∎
G.3 IPO do not satisfy the redescending property
The objective of IPO [37] is as follows:
$\widetilde{\mathcal{L}}_{\mathrm{IPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\coloneqq\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg[\bigg(\frac{g_{\theta}(\tilde{s})}{\beta}-\frac{1}{2\beta}\bigg)^{2}\bigg].$ (33)
We first show the IF for the IPO.
Theorem 8.
$\mathrm{IF}_{\mathrm{IPO}}(x,\theta,p_{\mathcal{D}})=-\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{IPO})}(s)\right]\bigg)^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{IPO})}(s_{\textrm{flip}})],$ (34)
where $F_{\theta^{*}}^{(\textrm{IPO})}(s_{\textrm{flip}})\coloneqq 2\bigg(\frac{g_{\theta}(s_{\textrm{flip}})}{\beta}-\frac{1}{2\beta}\bigg)\bigg(\nabla_{\theta}\log\pi_{\theta^{*}}(y_{\textrm{win}}^{\textrm{flip}}\mid x)-\nabla_{\theta}\log\pi_{\theta^{*}}(y_{\textrm{lose}}^{\textrm{flip}}\mid x)\bigg)$.
Proof.
$)].\displaystyle\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{IPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg[2\bigg(\frac{g_{\theta}(\tilde{s})}{\beta}-\frac{1}{2\beta}\bigg)\bigg(\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{win}}\mid\tilde{x})-\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{lose}}\mid\tilde{x}$$\bigg)\bigg$].
Corollary 5 (IPO is not robust).
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$, where $0<C<\infty$. Let $g_{\theta^{*}}(s)$ be bounded over $p_{\mathcal{D}}(s)$. Then, under Theorem 8, the IF of IPO do not satisfy the robustness condition in Definition 2, i.e., $\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{IPO}}(x,\theta,p_{\mathcal{D}})\|_{2}=\infty$.
Proof.
From Theorem 8, the IF for IPO is $\mathrm{IF}_{\mathrm{IPO}}=-(\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{IPO})}(s)])^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{IPO})}]$. From the positive definite assumption on the Hessian, let $L^{\prime}=\lambda_{\mathrm{min}}(\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{IPO})}(s)])>0$. The norm of its inverse is bounded: $\|(\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{IPO})}(s)])^{-1}\|\leq 1/L^{\prime}$. The gradient term $\|\nabla_{\theta}\log\pi_{\theta}(\dots)\|$ is also bounded by $2C$ from the assumption.
We analyze the limit of the IF:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{IPO}}\|$
$\leq(1/L^{\prime})\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{IPO})}(s_{\textrm{flip}})]\bigg\|$
$\leq(1/L^{\prime})\cdot\mathbb{E}_{p(s_{\textrm{flip}})}\bigg[\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\underbrace{2\cdot\bigg(\frac{g_{\theta}(s_{\textrm{flip}})}{\beta}-\frac{1}{2\beta}\bigg)}_{\text{IF Weight}}\cdot\underbrace{\bigg(\nabla_{\theta}\log\pi_{\theta^{*}}(y_{\textrm{win}}^{\textrm{flip}}\mid x)-\nabla_{\theta}\log\pi_{\theta^{*}}(y_{\textrm{lose}}^{\textrm{flip}}\mid x)\bigg)}_{\text{Gradient Term}}\bigg\|\bigg],$
where we use Fatou’s Lemma to exchange the limit.
In the limit, $\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}$, the IF weight term diverges:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\left\|2\cdot\left(\frac{g_{\theta}}{\beta}-\frac{1}{2\beta}\right)\right\|=\infty.$
Since the IF is proportional to the product of this diverging term ($\to\infty$) and a bounded, non-zero gradient term ($\leq 2C$), the IF itself diverges, that is,
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{IPO}}(x,\theta,p_{\mathcal{D}})\|=\infty.$
This completes the proof. ∎